Assume the car can be purchased for 0% down for 60 months (in lieu of rebate). A BMW that has a sticker price of $62,490 with factory and dealer rebates of $6,000 (a) Find the monthly payment if financed for 60 months at 0% APR. (Round your answer to the nearest cent.) $ (b) Find the monthly payment if financed at 2.5% add-on interest for 60 months. (Round your answer to the nearest cent.) $ (c) Use the APR approximation formula to find the APR for part (b). (Round your answer to one decimal place.) % (d) State whether the 0% APR or the 2.5% add-on rate should be preferred. 0% APR 2.5% add-on rate.
Answer #1

Part A

Monthly payment = 62490/60 = $1041.50

Part B

discount price = 62490-6000 = 56490

add-on-interest = 56490*2.5%*60/12 = 7061.25

total value = 56490+7061.25 = 63551.25

monthly payment = 63551.25/60 = $1059.18

Part C

APR = 2Nr/(N+1) = (2*60*2.5%)/(60+1) = 4.9%

Part D

0% APR should be preferred

Because monthly payment is lower in that case.